Two cyclists start biking from a trail’s start 3 hours apart. The second cyclist travels at 10 miles per hour and starts 3 hours after the first cyclist who is traveling at 6 miles per hour. How much time will pass before the second cyclist catches up with the first from the time the second cyclist started biking?
Answer by ankor@dixie-net.com(12692)

Two cyclists start biking from a trail’s start 3 hours apart. The second cyclist travels at 10 miles per hour and starts 3 hours after the first cyclist who is traveling at 6 miles per hour. How much time will pass before the second cyclist catches up with the first from the time the second cyclist started biking?
:
Let t = time the 2nd cyclist is on the road when he catches up
Then (t+3) = time the 1st cyclist is on the road.
:
When the 2nd cyclist catches up with the 1st, we know they will have traveled the same distance.
:
Write a distance equation: Distance = speed * time
:
10t = 6(t+3)
10t = 6t + 18
10t - 6t = 18
4t = 18
t = 18/4
t = 4.5 hrs (time of the 2nd cyclist:
:
Then the 1st cyclist time = 7.5 hrs
Check to see if both traveled the same distance
4.5(10) = 45 mi
7.5(6) = 45 mi also
:
Make sense to you, not that hard, right?